A large balloon contains 11.7 grams of Helium. What volume will the helium occupy at an altitude of 10,000.meters, where the atmospheric pressure is 0.262 atm and the temperature is -50.0 °C? *


1
Expert's answer
2021-04-09T04:05:10-0400

Molar mass of helium is 4.0g/ mol


Moles of Helium "= \frac{4}{11.7}=0.34 moles"


PV=nRT

0.262×V=0.34×8.3144598×223

0.262V=658.67

V=[658.67÷0.262]

Volume=2514